OpenStudy (anonymous):

suppose that sin alpha = -3/5 and cot Beta = -2/3. pi

OpenStudy (anonymous):

if \[\sin(\alpha)=-\frac{3}{5}\]then \[\cos(\alpha)=-\frac{4}{5}\] by the 2-4-5 right triangle in your head. \[\cos(2\alpha)=1-2\sin^2(\alpha)=1-2\times (\frac{3}{5})^2\]

OpenStudy (anonymous):

i get \[1-\frac{18}{25}=\frac{7}{25}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\cot(\beta)=-\frac{2}{3}\] means \[\sin(\beta)=-\frac{3}{\sqrt{13}}\] and \[\cos(\beta)=\frac{2}{\sqrt{5}}\]

OpenStudy (anonymous):

then use \[\sin(\alpha +\beta)=\sin(\alpha)\cos(\beta)+\sin(\beta)\cos(\alpha)\] by plugging in the numbers we found above

OpenStudy (anonymous):

i.e. \[-\frac{3}{5}\times\frac{2}{\sqrt{5}}+(-\frac{3}{\sqrt{13}})\times (-\frac{4}{5})\]